OpenStudy (anonymous):

Two angles are supplementary , one angle is 3 less than twice the other. find the dimensions of all angles

OpenStudy (anonymous):

x+y=180

OpenStudy (anonymous):

y=3-x

OpenStudy (anonymous):

is this part right?

OpenStudy (raden):

the second equation should be y = 2x - 3

OpenStudy (anonymous):

ok im stuck here

OpenStudy (anonymous):

x+(2x-3)=180

OpenStudy (anonymous):

3x+3=180?

OpenStudy (raden):

x+y=180 ... (1) y=2x-3 ... (2) subtitute the 1st equation too the 2nd x+2x-3=180 3x-3=180 add both sides by 3 3x-3+3 = 180+3 3x = 183

OpenStudy (anonymous):

61

OpenStudy (anonymous):

so the smaller angle would be 61 and larger 119?

OpenStudy (raden):

that's right then subtitute again the 2nd equation to get y

OpenStudy (raden):

yup

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

you're welcome :)